OpenStudy (anonymous):

A boat sails along the shore. To an observer, the boat appears to move at a speed of 22 m/s, and a man on the boat walking forward appears to have a speed of 23.1 m/s. How long does it take the man to move 3 m relative to the boat? A. 1.3 s B. 2.7 s C. 0.14 s D. 0.13 s

OpenStudy (matt101):

Relative to the observer, the boat is moving at 22 m/s while the man is moving at 23.1 m/s. That means the man, relative to the boat, is moving at 23.1-22 = 1.1 m/s. v=d/t, so t = d/v --> t = 3/1.1 = 2.7 s

OpenStudy (anonymous):

Thanks matt101